# Grounding-IQA: Multimodal Language Grounding Model for Image Quality Assessment

### v1

@@ 1 Introduction
Based on the task designed above, we construct a high-quality dataset, GIQA-160K, to enable existing MLLMs with grounding-IQA capabilities. This dataset comprises 160K instruction-tuning data with 40K images from diverse domains (e.g., artificial distortion and in-the-wild scenes). Specifically, the dataset corresponds to two sub-tasks: GIQA-DES includes 60K corresponding data, and GIQA-VQA contains 100K related data. To circumvent the time-consuming and costly process of manual annotation, we design an automated annotation pipeline. This system utilizes the public IQA dataset [42, 51] (with human-annotated description), to construct the corresponding dataset. (1) For GIQA-DES. The task includes detailed descriptions with coordinates. We generate the data through advanced vision [25] and language [6] models. Through these models, we extract and filter objects and corresponding coordinates from existing descriptions and images. Meanwhile, coordinates are expressed in natural language and attached to text. This avoids extra specialized tokens and ensures data compatibility. (2) For GIQA-VQA. Inspired by previous work [42, 50, 18], we construct the required data from the detailed descriptions in GIQA-DES via the LLM. We use specific QA templates (i.e., “Yes/No’, abbreviated as Y; “What/How/Why”, abbreviated as W) and emphasize location-specific objects to generate appropriate data. The coordinates are also combined with the generated QA.
@@ 3.2 Automated Annotation Pipeline
Figure 4: Compared with applying object name (“man”), utilizing the description phrase $\mathcal{T}_{q}$ (“the man wearing a white t-shirt”) can achieve more accurate detection results.
@@ 3.4 GIQA-Bench
(3) Grounding precision. Measure the grounding performance for both GIQA-DES and GIQA-VQA. We use category-agnostic Intersection over Union (IoU) to evaluate box quality. We also define Tag-Recall to assess category-specific grounding capabilities. In Tag-Recall, a result is true positive (TP) only if both the IoU and object name similarity exceeds a 0.5 threshold. For fairness, the bounding box is represented by the normalized corner coordinate.
@@ 4.2 Ablation Study
Box Optimization. We evaluate box optimization in annotation pipeline, including box refinement (filter and merge) and representation. We compare the models trained on GIQA-DES with (Ref-Box) and without refinement (Raw-Box) in Tab. LABEL:tab:ablation_box_1. The refinement enhances the fine-tuning effect. We also visualize box area distribution in Fig. 6. Refinement reduces the difference between automatically annotated GIQA-160K and human-annotated GIQA-Bench.
@@ 4.3 Results on GIQA-Bench
In GIQA-Bench, we compare four groups of MLLMs with different functionalities, i.e., (1) General models (General): LLaVA-v1.5-7B [23], LLaVA-v1.5-13B [23], LLaVA-v1.6-7B [24], and mPLUG-Owl2-7B [48]; (2) Multimodal referring and grounding models (Ground): Shikra-7B [4], Kosmos-2-1.6B [34], Ferret-7B [50], and GroundingGPT-7B [18]; (3) IQA models (IQA): DepictQA-Wild-7B [51] and Q-Instruct [42] (fine-tuned three base models); and (4) Our methods (Ours): Four general models fine-tuned on GIQA-160K. Detailed testing prompts for all models are provided in the supplementary material.
Quantitative Results. We evaluate all models on GIQA-DES and GIQA-VQA tasks from two aspects: quality assessment and referring/grounding ability, as in Tab. 5. General models perform poorly on both tasks, while task-specific models are more effective in their respective domains. Specifically, grounding MLLMs (e.g., Shikra [4]) perform well on general grounding tasks, but show decreased performance when dealing with quality-related objects/areas (GIQA-VQA, Tag-Recall). Conversely, IQA models (e.g., Q-Instruct [42]) excel in description quality (GIQA-DES, LLM-Score), but exhibit low accuracy in GIQA-VQA. In contrast, our models outperform MLLMs specialized in grounding or IQA tasks in both aspects.
Qualitative Results. We provide some visual comparisons in Fig. 7. For GIQA-DES (left instance), the quality descriptions generated by general (mPLUG-Owl2-7B [48]) and grounding (Ferret [50]) MLLMs are unsatisfactory. In addition, the ground results of Ferret include many quality irrelevant objects, which is impractical for real applications. In contrast, our method describes image quality more properly with coordinates of key objects affecting the quality. Furthermore, in the GIQA-VQA task (right instance), our method produces more accurate responses to image quality VQA involving spatial perception. More qualitative results are provided in the supplementary material.

### v2

@@ 1 Introduction
Since existing datasets can not realize grounding-IQA well [21, 51, 43], we construct a new dataset, GIQA-160K, based on the proposed paradigm. This dataset can enhance the grounding-IQA capabilities of current MLLMs. The dataset comprises 160K instruction-tuning data with 40K images from diverse domains. Specifically, the dataset corresponds to two sub-tasks: GIQA-DES includes 60K corresponding data, and GIQA-VQA contains 100K related data. To construct the corresponding dataset, we design an automated annotation pipeline. The automated pipeline generates the GIQA-160K through the public IQA dataset [43, 52] (with the human-annotated description). (1) For GIQA-DES. The task includes detailed descriptions with coordinates. We generate the data through advanced vision [24] and language [6] models. Through these models, we extract and filter objects and corresponding coordinates from existing descriptions and images. Meanwhile, coordinates are expressed in natural language and attached to text. This avoids extra specialized tokens and ensures data compatibility. (2) For GIQA-VQA. Inspired by previous work [43, 51, 18], we construct the required data from the detailed descriptions in GIQA-DES via the LLM. We use specific QA templates (i.e., “Yes/No”, abbreviated as Y; “What/How/Why”, abbreviated as W) and emphasize location-specific objects to generate appropriate data. The coordinates are also combined with the generated QA.
@@ 3.2 Automated Annotation Pipeline
Figure 4: Compared with applying object name (“man”), utilizing the description phrase $\mathcal{T}_{r}$ (“the man wearing a white t-shirt”) can achieve more accurate detection results.
@@ 3.4 GIQA-Bench
(3) Grounding Precision. Measure the grounding performance for both GIQA-DES and GIQA-VQA. We use category-agnostic mean Intersection over Union (mIoU) to evaluate box quality. We also define Tag-Recall to assess category-specific grounding capabilities. In Tag-Recall, a result is true positive only if both the IoU and object name similarity exceeds a 0.5 threshold. For fairness, the bounding box is represented by the normalized corner coordinate.
@@ 4.2 Ablation Study
Box Optimization. We evaluate box optimization in annotation pipeline, including box refinement (filter and merge) and representation. We compare the models trained on GIQA-DES with (Ref-Box) and without refinement (Raw-Box) in Tab. LABEL:tab:ablation_box_1. The refinement enhances the fine-tuning effect. We also visualize box area distribution in Fig. 6. Refinement reduces the difference between automatically annotated GIQA-160K and human-annotated GIQA-Bench. Besides, more analyses are provided in the supplement.
@@ 4.3 Results on GIQA-Bench
In GIQA-Bench, we compare four groups of MLLMs with different functionalities, i.e., (1) General models (General): LLaVA-v1.5-7B [22], LLaVA-v1.5-13B [22], LLaVA-v1.6-7B [23], and mPLUG-Owl2-7B [49]; (2) Multimodal referring and grounding models (Ground): Shikra-7B [4], Kosmos-2-1.6B [33], Ferret-7B [51], and GroundingGPT-7B [18]; (3) IQA models (IQA): DepictQA-Wild-7B [52] and Q-Instruct [43] (fine-tuned three base models); and (4) Our methods (Ours): Four general models fine-tuned on GIQA-160K. The detailed test settings and analyses are provided in the supplementary material.
Quantitative Results. We evaluate all models on GIQA-DES and GIQA-VQA tasks from two aspects: quality assessment and referring/grounding ability, as in Tab. 5. General models perform poorly on both tasks, while task-specific models are more effective in their respective domains. Specifically, grounding MLLMs excel in general grounding tasks but underperform on quality-related objects/areas (GIQA-VQA, Tag-Recall). Conversely, IQA models achieve high description quality (GIQA-DES, LLM-Score), but exhibit low accuracy in GIQA-VQA. In contrast, our method outperforms existing MLLMs. Moreover, to further reveal the IQA capability of our method, we evaluate it on traditional score-based IQA benchmarks. Our method also achieves impressive performance. Details are provided in the supplementary material.
Qualitative Results. We provide some visual comparisons in Fig. 7. For GIQA-DES (left instance), the quality descriptions generated by general (mPLUG-Owl2-7B [49]) and grounding (Ferret [51]) MLLMs are unsatisfactory. In contrast, our method describes image quality more properly with coordinates of key objects affecting the quality. Furthermore, in the GIQA-VQA task (right instance), our method produces more accurate responses to image quality VQA involving spatial perception. More qualitative results are provided in the supplementary material.